Helllllp I don't know how to find a discriminant. u_u
ex: 2x^2+5x+6=0


d=
# of real number solutions:

Also, suppose a parabola has a vertex in Quadrant I and a<0 in the equation y=ax^2+bx+c. How many real solutions will the equation ax^2+bx+c=o have?

Answer by lawdej(58)


here is the way i hope you will be satisfied
for a quadratic equation ax^2+bx+c=0 the discriminant is b^2-4ac
so here a=2,b=5,c=6
thus b^2-4ac=5^2-(4x2x6)=25-48=-23 this indicates the equation has no solution
in general if b^2-4ac<0 no solution,if b^2-4ac=0 one solution,if b^2-4ac>0 two solution